OpenStudy (anonymous):

Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

i would use elimination because if you add the equations the \(y\) term drops out in one step

OpenStudy (thesecret20111):

I would use elimination as look at the terms in each equation, adding them together would cancel out y. This would make solving for x easy. Once you have x you could then solve for y.

OpenStudy (anonymous):

Can you help me solve for it please?

OpenStudy (anonymous):

add

OpenStudy (thesecret20111):

Ok so as I said add both equations together.. What would this give you?

OpenStudy (anonymous):

\[ 2x + 5y = −3\]\[ −15x − 5y = −10 \]

OpenStudy (anonymous):

i.e. add in columns

OpenStudy (thesecret20111):

So add the 2x with the -15x and the +5x with the -5y and the -3 with the -10

OpenStudy (anonymous):

2x+5y=-3 -15x-5y=-10 Cross out the y's. 2x-15x=-10 13x=-10-3 13x=-13 Divide by 13 x=1 2(1)+5y=-3 2+5y=-3 -2 -2 5y=-5 Divide by 5 y=-1 (1,-1)

OpenStudy (anonymous):

your answer is right but your method has a flaw

OpenStudy (anonymous):

you should have in line 3 \[-13x=-13\] so \(x=1\)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw